ANALYTIC HYPOELLIPTICITY IN DIMENSION TWO

MICHAEL CHRIST

Abstract. A simple geometric condition is sufficient for analytic hypoellipticity of sums of squares of two vector fields in ℝ^{2}. This condition is proved to be necessary for generic vector fields and for various special cases, and to be both necessary and sufficient for a closely related family of operators.